### v1

# Associative Knowledge Graphs for Efficient Sequence Storage and Retrieval

Abstract
This paper presents a novel approach for constructing associative knowledge graphs that are highly effective for storing and recognizing sequences. The graph is created by representing overlapping sequences of objects, as tightly connected clusters within the larger graph. Individual objects (represented as nodes) can be a part of multiple sequences or appear repeatedly within a single sequence. To retrieve sequences, we leverage context, providing a subset of objects that triggers an association with the complete sequence. The system’s memory capacity is determined by the size of the graph and the density of its connections. We have theoretically derived the relationships between the critical density of the graph and the memory capacity for storing sequences. The critical density is the point beyond which error-free sequence reconstruction becomes impossible. Furthermore, we have developed an efficient algorithm for ordering elements within a sequence. Through extensive experiments with various types of sequences, we have confirmed the validity of these relationships. This approach has potential applications in diverse fields, such as anomaly detection in financial transactions or predicting user behavior based on past actions.
Keywords structural associative knowledge graphs  $\cdot$ graph density  $\cdot$ sequence storage  $\cdot$ sequence retrieval  $\cdot$ context-based association
1 Introduction
[…]
Transformers [3], introduced later, eliminate the need for recurrent and convolutional neural networks, offering faster and more parallelizable structures. Hierarchical Temporal Memory (HTM), by [4] draws inspiration from the human brain’s structure. Projects like nupic.torch [5] integrate HTM principles into PyTorch highlighting the benefits of sparse representations.
Associative Long Short Term Memory LSTM [6] significantly improves d the performance of LSTM in tasks requiring storing and retrieving complex data by adding special associative memory. Clustered Neural Network (CNN) has the ability to recall stored messages quickly and enables full parallel processing, and its modification Restricted Clustered Neural Network (R-CNN) [7], [8] allows to reproduce data with fewer errors.
Knowledge Graphs offer a powerful framework for representing objects and their relationships, enabling complex decision-making [9]. Graph Neural Networks (GNNs [10]) has further enhanced the ability to process graph-structured data using deep learning techniques. The ANAKG (Active Neuro-Associative Knowledge Graphs [11], [12]) utilize dynamic neurons for learning sequences within an associative knowledge graph.
Our Contribution:
Our previous work on structural associative knowledge graphs [13] demonstrated the potential of this approach for storing and retrieving scenes. In this paper, we extend this research to sequence storage and retrieval.
We show that the same structural principles can be effectively applied to sequences, preserving the key properties of memory size, graph density, and context requirements.
A key challenge in applying structural approaches to sequence storage is the efficient ordering of retrieved sequence elements. Our key contribution lies in developing efficient algorithms for ordering retrieved sequence elements. This paper compares the efficiency of different sequence ordering algorithms and presents testing results on various datasets.
The program’s code is available on GitHub [14].
2 Building a Knowledge Graph
Associative knowledge graphs were developed to rapidly construct semantic memories that store associations between observed events, actions, and objects. These graphs form the foundation of knowledge about the outside world, making them invaluable for autonomous learning systems in open environments.
In this paper, we use structural associative knowledge graphs to store and retrieve sequences involving diverse objects. In these graphs, the structure itself is used to recreate the contents of memory based on a given context, and interconnection weights are then used to restore the proper order of the stored sequence elements.
The structure of the knowledge graph develops automatically by progressively adding information, such as sentences or sequences of numbers. Synaptic connections and their strengths between represented elements are recorded during the graph’s creation. Each node in the knowledge graph represents a specific object, word, or concept. As knowledge is input, these nodes become interconnected. However, when a large amount of information is recorded with a fixed number of nodes, synaptic connections may undergo multiple modifications, reducing or losing the resolution of the original information.
In this work, we extensively analyze the structural properties of knowledge graphs, emphasizing the importance of knowledge graph density, measured by the ratio of used synaptic connections to the total possible synaptic connections among the nodes. We show that the theoretical analysis performed for scene memory [13] can be applied to sequence memory. Based on this analysis, we accurately estimate the most important dependencies between the number of nodes in the graph, graph density, context size, and the number of saved sequences. This allows us to determine and use the appropriate graph size or necessary context based on the size and number of sequences stored within the associative knowledge graph memory (referred to as graph memory).
A crucial aspect of reproducing a sequence is the proper ordering of its elements. An added challenge is that the given context does not necessarily preserve the order of the context elements in the stored sequence. A key contribution of this paper is demonstrating that the sparsity of the graph memory enables high memory capacity, which increases quadratically with the size of the graph.
Definition 1.
Sequential Structural Associative Knowledge Graph (SSAKG) is a graph in which each sequence is represented by a corresponding transitive tournament.
The SSAKG (referred to as a sequential graph or a graph) density becomes a crucial factor influencing memory storage capacity. However, the sequences stored should be sufficiently intricate to enable sequence retrieval with a relatively small context (i.e., the number of observed objects of a sequence). Additionally, to maintain graph sparsity, a large number of unique objects must be used in the stored sequences.
Sequences loaded into memory are represented by the corresponding directed graph in which each vertex is connected to all vertices with higher numbers than itself. Such graph is called a transitive tournament. The transitive tournament is characterized by its hierarchical ordering and its acyclic nature where there is a single directed edge between every pair of vertices.
Each transitive tournament will be placed in the memory graph (SSAKG). The individual elements of the sequence will be placed in the vertices of the graph with corresponding numbers. The order of the elements corresponds to the corresponding connections, i.e. the first element connects to all the others, the second to all the others except the first, and so on.
Example 1.
As an example, let’s consider a SSAKG graph that contains 20 vertices and can store sequences consisting of numbers 1 to 20. Let’s assume that we first store the sequence [2,6,11] in this graph. The first element of this sequence (no. 2) is connected via the directed edges to the numbers 6 and 11. The second element (no. 6) connects to no. 11. No edge is derived from element no. 11 (Figure 1a). Figure 1b) shows the recorded sequence [2,6,11] and the sequence [11,8,2].
Figure 1: Sequences stored in the memory graph. Figure a) shows the stored sequence [2,6,11]. Figure b) shows the stored sequences [2,6,11] and [11,8,2].
Definition 2.
Sequential graph memory storage capacity is the maximum number of sequences that can be uniquely recalled without errors.
Both aspects of this definition are essential. A recalled sequence is considered error-free if all elements of the sequence are correctly recalled. A recalled sequence is deemed unique if no other sequence matching the provided context is recalled.
In the proposed approach to associative memory, where transitive tournament subgraphs represent individual sequences, the storage capacity strongly depends on the graph’s density. This density is influenced by the number of nodes (representing objects) and the size and quantity of stored sequences. As more sequences are added, the graph’s density increases. However, once the density surpasses a critical threshold, recalling stored memories without error becomes impossible. Therefore, determining this critical density is crucial as it defines the memory capacity.
When a knowledge graph becomes overly dense due to a limited number of unique objects, virtual objects can be a valuable solution. Virtual objects help maintain a sparse graph structure even when dealing with a large number of data. They are created by combining an object’s symbol with its location in the sequence. This limitation, the new approach presented in this work is necessary.
To further differentiate virtual objects, additional characteristics like color, position, or size can be incorporated. While it’s theoretically possible to reconstruct a sequence using the scene memory approach described in [13] if all subsequent elements are converted to virtual objects, this would require context elements to also include their precise sequence location. To avoid this limitation, the new approach presented in this work is necessary.
This section investigates the relationship between the density of a SSAKG graph and the quantity of stored sequences. As the number of sequences grows, so does the graph density influenced by factors such as average sequence length and the total number of nodes. The overlapping connections between densely connected clusters (known as transitive tournament subgraphs) contribute to the overall graph density. To address this overlap and maintain efficiency, we must consider the effective number of synaptic connections within a SSAKG that stores multiple transitive tournament subgraphs.
If the density $d_{i}$ of a directed graph spanning $n$ vertices is known, then the number of oriented edges in this graph can be calculated from the formula
$e_{i}=d_{i}\ast n\ast\left(n-1\right)$ (1)
By incorporating additional transitive tournament to the existing graph, we add $e_{a}$ additional edges
$e_{a}=(1-d_{i})\ast l_{s}\ast\left(l_{s}-1\right)$ (2)
where $l_{s}$ is the length of the added sequence.
Therefore, each new sequence written in the graph increases the graph density according to the relationship
$d_{i+1}=\frac{e_{i}+e_{a}}{e_{n}}=\frac{d_{i}\ast n\ast\left(n-1\right)+\left(1-d_{i}\right)\ast l_{s}\ast\left(l_{s}-1\right)}{n\ast\left(n-1\right)}$ (3)
Where $e_{n}$ represents the maximum number of oriented connections in the knowledge graph spanned on $n$ nodes. After simplification, we have
$d_{i+1}=d_{i}\ast\left(1-\frac{l_{s}\ast\left(l_{s}-1\right)}{n\ast\left(n-1\right)}\right)+\frac{l_{s}\ast\left(l_{s}-1\right)}{n\ast\left(n-1\right)}=d_{i}\ast\left(1-\xi\right)+\xi$ (4)
where
$\xi=\frac{l_{s}\ast\left(l_{s}-1\right)}{n\ast\left(n-1\right)}$ (5)
We can now easily connect the final density of the graph with the number of sequences $s$ stored in it. Since
$d_{0}=0$ (6)
we have the final graph density
$d_{s}=1-\left(1-\xi\right)^{s}$ (7)
and
$s=\frac{\log\left(1-d\right)}{\log\left(1-\xi\right)}.$ (8)
[…]
3 Creating a Structured Sequence Memory
To establish a sequence memory, we can adapt the existing method used for structural scene memory. The key difference lies in the representation: instead of undirected complete graphs, represented by a full square matrix, scenes are now articulated as upper triangular matrices. This results in an asymmetric matrix that signifies a directed knowledge graph.
To reconstruct sequences based on a given context, we can follow a similar approach to recreating scenes in structured memories. This involves adding an asymmetric knowledge graph matrix and its transpose to create a symmetric matrix. Within this matrix, sequence elements (vertices) are selected in a manner analogous to finding scene elements based on context.
Ultimately, the maximum density of a knowledge graph representing sequential memory will be similar to that of a scene memory.
In a directed graph, multiple sequences can be stored and retrieved based on context. As more sequences are added, the graph’s density increases. We can estimate this density using a method similar to the one used for structured scene memories [13].
It is important to note that adding a directed graph matrix and its transpose can potentially double the density of the resulting symmetric matrix. Therefore, the maximum density of the knowledge graph used for recording sequences might be limited to half that of scene memory. However, this effect is offset by the fact that triangular matrices are sufficient for recording sequences, while square matrices with nearly twice as many non-zero elements are needed for scenes. Ultimately, the maximum density of a knowledge graph representing sequential memory will be similar to that of scene memory.
4 Sequence Elements Ordering Algorithms
To recover a sequence from an associative knowledge graph the retrieved elements must be ordered. This is particularly important for large datasets, as it allows us to convert the retrieved data into a properly ordered "transitive tournament graph matrix." While a simple sorting method might suffice for small, sparse graphs, more sophisticated techniques are needed for denser graphs containing elements from multiple sequences.
This section explores four algorithms for sequence element ordering:
  1. 1.
    Simple Sort
  2. 2.
    Node Ordering
  3. 3.
    Enhanced Node Ordering
  4. 4.
    Weighted Edges Node Ordering.
Simple Sort.
This baseline algorithm sorts the retrieved data (represented by rows in matrix M) based on the number of non-zero elements. It does not remove elements that do not belong to the target sequence and serves as a reference point for comparison.
Node Ordering.
This algorithm aims to generate all possible orderings of rows and columns in matrix M to obtain an upper triangular matrix (without a diagonal) containing only ones in the upper right corner. It prioritizes rows with n-1 non-zero elements and removes both the identified row and its corresponding column. This process is repeated for all such rows. If rows have the same number of non-zero elements, all possible ordering combinations are considered. A key difference from Simple Sort is the removal of additional elements from potentially different sequences. For instance, when selecting the first row with n-1 non-zero elements, the entire corresponding column is removed, potentially containing elements from other sequences. Figure 2 illustrates the algorithm’s block diagram. To illustrate the results of the Node Ordering algorithm, let’s examine a straightforward example.
Example 2.
Consider $55x$ 5 matrix M and assume that the path obtained by the Node Ordering algorithm consists of indices $[4,3,3,2,1]$. These indices represent the progressively eliminated rows (starting from row 4) of successively smaller matrices. Initially, with the original numbering $[1,2,3,4,5]$, we arrive at smaller and smaller matrices, which align with the original row numbers as illustrated here: $[1,2,3,5]\to[1,2,5]\to[1,2]\to[1]$. This path corresponds to the desired sequence elements renumbering: $[4,3,5,2,1]$.
It’s important to note that this procedure is reversible because, given the renumbered rows, we can determine the path generated by the Node Ordering algorithm. The block diagram of all the Node Ordering algorithms is shown in Fig.2 a), and Fig. 2 b) shows the structure of Prioritize functions that are different for each algorithm.
Figure 2: a) Block diagram of the Node Ordering. b) Prioritize functions that are different for each algorithm.
Enhanced Node Ordering.
This approach builds upon the Node Ordering Algorithm by considering the weights associated with connections between individual vertices of the directed graph. Ideally, the Node Ordering algorithm should identify only one row with n-1 number of non-zero elements. However, as the density of the graph increases, more and more additional elements from other sequences appear in the rows, making it more likely to find rows with the same number of non-zero elements. In such cases, the algorithm prioritizes edges with higher weights, reducing the number of alternative orderings.
Weighted Edges Node Ordering.
This method leverages weights directly embedded in the transitive tournament graph matrix (presented in Example 1). By multiplying the individual rows of the graph matrix by successive numbers, the algorithm can find rows with appropriate weight values.
5 Comparison of Graph Ordering Algorithms Performance
Figure 3 compares the performance of all four algorithms in reconstructing sequences. The test involved 1000 sequences, each containing 15 elements. Figure 4a shows the results for a graph containing 1000 elements, while Figure 4b presents the results for a 2000-element graph.
In both cases, the context used for sequence reconstruction comprised 7 elements. Histograms illustrate the reproducibility of sequence elements. The x-axis represents the number of correctly reproduced elements, with 15 indicating perfect reconstruction. As evident, the Weighted Edges Ordering algorithm outperforms all others.
Figure 3: Different SSAGK graph sizes used based on a) 1000 elements, b) 2000 elements. The 1000 sequences were memorized. As you add repeated sequence elements, the size of the graph matrix increases.
Figure 4 explores the average number of node orderings obtained for sequences stored in SSAKG graphs of varying sizes (all sequences were 10 elements long) under the assumption that all sequence elements are known.
Figure 4: Comparison of path sorting algorithms. Different graph memory (SSAKG) dimension was used a) 500, b) 2500. The sequence length is 10. The image shows results averaged over 1,000 different memory graphs.
As expected, the average number of orderings decreases with increasing graph size (represented on the lower x-axis) and increasing graph density (upper x-axis). This experiment focuses on Algorithms 2 (Node Ordering) and 4 (Weighted Edges Node Ordering) since Simple Sort performed significantly worse, and Enhanced Node Ordering offered similar results to Node Ordering. The figure clearly demonstrates that the Weighted Edges Node Ordering algorithm exhibits a much slower growth in the average number of node orderings, indicating its superior efficiency in accurately reconstructing sequences stored in the SSAKG graph.
Overall, these results highlight the importance of effective sequence node ordering algorithms for accurate data retrieval from associative knowledge graphs. The Weighted Edges Node Ordering algorithm demonstrates superior performance compared to simpler approaches.
6 Practical Examples
This section presents examples based on memorizing and recalling sentences from SSAKG memory. To obtain correct sentences, the program uses the NLTK library - Natural Language Toolkit. [15]. This library enables the reading of sentences and words with appropriate grammar for the specified language. Additionally, it contains a vast collection of texts grouped into “corpora". For our experiments, we utilized a corpus based on Gutenberg’s project [16], which includes several selected book entries. We carefully selected several thousand sentences with interesting parameters, filtering out common words (stop words) and punctuation marks [17].
Example 3.
Sentences of Equal Length (15 words). We selected 1000 random sentences, each containing 15 words. These sentences were prepared using the previously described method. The random text contained 4453 different words. The word distributions in the text are depicted in Figure 5. The selected sentences were then encoded into SSAKG.
Figure 5: Words occurring in the text a) distribution of all words in the text, b) distribution of the most frequent words.
Subsequently, using contexts of length 8, 9, and 10, we attempted to recall these sentences from memory. The results are presented in Table 1. For randomly generated sequences with a flat distribution, the program successfully reproduced all sequences.
Table 1: Correctness of Reading Sentences with 15 Words
Context Length Correct Sets of Words Correct Order
8 95.1% 96.3%
9 96.6% 96.1%
10 97.3% 95.9%
Example 4.
Sentences of Varying Lengths (10-15 words). We selected 1000 random sentences with lengths ranging from 10 to 15 words. The sentences were prepared using the previously described method. This random text contained 3037 different words. The word distributions in the text are depicted in Figure 6. The selected sentences are then encoded into SSAKG. Subsequently, using contexts of length 8,9, and 10, we attempted to recall these sentences from memory.
Figure 6: Words occurring in the text a) distribution of all words in the text, b) distribution of the most frequent words.
The results are presented in Table 2. For randomly generated sequences with a flat distribution, the program successfully reproduced all sequences.
Table 2: Correctness of Reading Sentences with 10-15 Words
Context Length Correct Sets of Words Correct Order
8 94.7% 95.4%
9 95.2% 94.7%
10 97.0% 94.7%
7 Conclusions
This paper introduces a novel structural approach for storing and retrieving sequences based on a given context. We employed associative knowledge graphs composed of transitive tournaments that represent stored sequences. Our research demonstrates that these graphs can effectively create associative sequential memories by utilizing structural information about synaptic connections and integer weights representing sequence element order.
We found that the memory capacity is determined by the size of the knowledge graph and the density of its synaptic connections. Our theoretical analysis reveals that memory capacity grows quadratically with the number of neurons used to build the knowledge graph. While smaller graphs may not yield significant sequence capacity, larger graphs can store a considerable number of sequences, allowing each neuron to represent multiple sequences.
A crucial aspect of our work was the development of algorithms for ordering sequence elements from the retrieved set of graph nodes activated by the input context. This paper presents results comparing the performance of these algorithms.
We conducted memory tests using both randomly generated synthetic data and real-world datasets, such as a collection of sequences of words featuring different sentences. These tests confirmed the validity of our findings.
Our approach offers a novel way to store sequences efficiently. By leveraging the sparse nature of knowledge graphs and the relatively small size of recorded sequences, we can achieve significant memory capacity for sequences. Unlike other neural networks used for sequence storage, our method has several advantages:
  • No training is required: The associative graph structure is created automatically as new sequences are added.
  • Flexible context: The context used to search for and reconstruct sequences can be presented in any order.
  • Efficient retrieval: Even a relatively small context is sufficient to reconstruct the entire sequence.
Our future work related to this topic will focus on two key aspects. First, we aim to compare the proposed structural associative memory with contemporary associative memory solutions, such as LSTM or NuPick. A comparative analysis of these two solutions will be the subject of our upcoming article. Second, we plan to enhance the presented node ordering concept by incorporating novel elements like using prime numbers to represent sequence order and micro-columns to increase memory capacity.